OpenStudy (anonymous):

What is the slope of the line passing through points A (5,4) and B(0,3)?

OpenStudy (amistre64):

subtract one point from the other and stack the y over x to get slope

OpenStudy (amistre64):

x,y (5,4) -(0,3) ------ 5,1; slope = y/x = 1/5

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

amistre64 you are awsome!

OpenStudy (amistre64):

lol .. i can be :)

OpenStudy (anonymous):

always remember slope is change in y over change in x. With 2 points you can get the change by subtracting the coordinates: (y1 - y0)/(x1 - x0).

OpenStudy (anonymous):

slope= y1-y2/x1-x2

OpenStudy (anonymous):

slope suppose if m then , \[m =\left( y1-y0 \right) \div \left( x1-x0 \right)\]